OpenStudy (anonymous):

HELP. use substitution to solve each system of equations: 1. 1) Y=x+5 2) 4x+y=2

OpenStudy (anonymous):

x=y-5 4(y-5)+y=2 4y-20+y=2 5y-20=2 5y=22 y=22/5 x=22/5-5 x=-3/5, y=22/5

OpenStudy (jonnyvonny):

Do you require further elaboration, lashae?

OpenStudy (anonymous):

how did you get -5

OpenStudy (jonnyvonny):

He messed up, is all.